The p-system with Large Data

Robin Young Univ of Massachusetts

We describe recent progress towards proving global existence of solutions to the p-system of isentropic gas dynamics. The Cauchy data may be large in amplitude and in total variation. We give precise conditions for the formation and annihilation of a vacuum. We then describe a modified front tracking scheme for constructing large amplitude solutions.